OpenStudy (anonymous):

Find the soulution if possible 2(x+2)-3=2x+1

OpenStudy (anonymous):

2(x+2) - 3 = 2x + 1 this is given now see 2x + 4 - 3 = 2x + 1 2x + 1 = 2x + 1 transpose 1 to right hand side 2x = 2x + 1 - 1 2x = 2x haha it proved that 1 = 1 sorry no solution is possible if the above possibilities are given. no from my side :)

OpenStudy (anonymous):

2x + 4 - 3 = 2x + 1 2x + 1 = 2x + 1 therefore all numbers work in this equation

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

yes i think no solution because 2x + 1 itself is the simplified form of 2(x + 2) - 3

OpenStudy (anonymous):

(-infinity to plus infinity :P)

OpenStudy (anonymous):

so when its no soulution its infinitely?

OpenStudy (anonymous):

infinity?

OpenStudy (anonymous):

i don't think so.. the answer should be NO SOLUTION

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

hey when u can answer it its the answer and therefore there is a souloution correct>?

OpenStudy (anonymous):

if you get one solution for the first degree.. then yes a solution exists.. similary for a quadratic you have to get 2 answers, and cubic 3.. and so on!

OpenStudy (anonymous):

thanx have a great day